OpenStudy (anonymous):

Solve the following quadratic equation. 3(x + 5)^2 = 60

OpenStudy (anonymous):

Solve the following quadratic equation. 3(x + 5)2^ = 60

OpenStudy (anonymous):

Use the quadratic formula to solve the following equation. 3a - 4 = a(a - 3)